OpenStudy (anonymous):

write 99 as a product of primes.

OpenStudy (anonymous):

Prime Factorize 99

OpenStudy (raden):

99 = 9 * 11 = 3^2 * 11

OpenStudy (anonymous):

99 = 9 x 11 = 3^2 * 11

OpenStudy (anonymous):